OpenStudy (anonymous):

6 3/5 - 2 1/4 @dakid88

OpenStudy (dakid88):

8 and 17/20

OpenStudy (anonymous):

hes wrong

OpenStudy (anonymous):

what is it then?

OpenStudy (anonymous):

4and 7/20

OpenStudy (anonymous):

you should turn them into improper fractions then find a number that 5 and 4 go into and then subtract the numerators not the denominators, then simplify

OpenStudy (anonymous):

thats simplified

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

so 6 3/5=33/5

OpenStudy (dakid88):

how am im wrong

OpenStudy (anonymous):

thank you @ash_heather1

OpenStudy (anonymous):

he said you didn't simplify @dakid88

OpenStudy (anonymous):

and 2 1/4= 9/4

OpenStudy (anonymous):

ok^

OpenStudy (dakid88):

oh yeah i forgot

OpenStudy (anonymous):

its okay

OpenStudy (anonymous):

20 goes into 5&4 so multiply the numerators how many times you multiplied the denominators to get to 20

OpenStudy (dakid88):

wait u cant simplyfy 8 and 17/20

OpenStudy (anonymous):

ok @ash_heather1

OpenStudy (anonymous):

hey sorry i was away

OpenStudy (anonymous):

so 9/4 would be 45/20 because 4 goes into 20 5 times so u multiply 9*5 and put it over 20

OpenStudy (anonymous):

and 33/5 would be 132/20

OpenStudy (anonymous):

then subtract the numerators and keep the denominators the same so 132-45=87, and then have 87 over 20 so 87/20 simplified is 4 7/20